Pongtronics Pongtronics
- Unit Log -

1976: Constructed.
1997~: Refurbished. U1 (oscillators) replaced with 4011. Should be a 4001, but doesn't matter in the U1 application.
2008 Dec: Intermittent operation of paddles. Bad contacts on Cinch connectors. Improved with contact cleaner.
2008 Dec: Loss of borders and most video. U9 failed (Fairchild 4001). Replaced with MC4001B.
2008 Dec: D flip-flop (U31) added to synchronise top and bottom borders to horizontal scan. This was not provided for in the original design and resulted in a jitter on the borders as the borders could begin and end at an arbitrary point along a horizontal scan line. The addition ensures the borders are composed of complete horizontal lines.
2009 Feb 07: Display shows tearing of raster at top of screen. V-sync pulse width is too long, around 1.5mS. 10K resistor in V oscillator replaced with 2.49K, bringing the pulse width down to 0.44mS.
2009 Feb 09: V-sync pulse still too long by RS-170 spec. V oscillator resistor replaced again, with 1K, bringing the pulse width down to 0.2mS. 100K resistor replaced with 130K. 10K resistor in V-ramp integrator replaced with 1K, to bring ramp reset time within V-sync pulse.
2009 Feb 13: New sync generation circuits tried out. These circuits derive both horizontal and vertical sync from a single oscillator, so the phase relationship between the two is constant. This significantly reduces jitter in the display.
2009 Feb 13: Left score bar noticed to be incrementing multiple units when external sync oscillator is present. Glitches on low end of Ball H integrator output at end of V-sync pulse result in multiple triggerring of the left score counter. Not clear where glitches are coming from. 1M resistor in ground level 'miss' comparator reduced to 560K to increase hysteresis. May have to be reduced further, but behaviour is idiosyncratic.
2009 Feb 19: 4011 in U1 position replaced with 4001.
2013 Feb: Poor H sync. H oscillator frequency down around 14 KHz. Adjusted.
2013 Feb: Paddles jumpy. Contact cleaner applied to paddle pots. Much improved.
2013 Feb: No score bars present. No outputs from U24 (4024). Replaced.

Display observed to be better without score bars. Score bars seem to upset V sync. High frequency oscillation noted on 'scope on edges of score bar video. Probably due to analog input with poor slopes on CMOS gate inputs. Should try to improve. Video signal levels and noise should be examined for improvement.

2008 Dec